OpenStudy (anonymous):

Find the domain of the function: g(x)=√7-x

Parth (parthkohli):

\[\sqrt{7 - x }\ge 0\]Does that make sense?

OpenStudy (lgbasallote):

actually..the solution should be \[7 - x \ge 0\] no square root symbols

Parth (parthkohli):

That'd be the same thing, because you get @lgbasallote's equation after squaring both sides.

OpenStudy (lgbasallote):

algebraically they're equivalent..logically no

OpenStudy (anonymous):

Thanks! So it is greater than or equal to because you can't take the square root of a negative, right?

OpenStudy (anonymous):